OpenStudy (okdutchman7):

Please help. WILL MEDAL! @jim_thompson5910 @Mehek14 @hartnn @johnweldon1993 @satellite73 @dumbcow

OpenStudy (okdutchman7):

Determine if \[f(x) = \frac{ 1 }{ x-3 }\] and \[g(x) = \frac{ 3x+1 }{ x }\] are inverse fractions by computing their compositions.

OpenStudy (ask_if_im_a_ghost):

By definition: f(x) and g(x) are inverses of each other if f(g(x))=x or g(f(x))=x So simply compute g(f(x)) and if it equals x, then they are inverse.